You wish to invest 2,000 pesos in Uruguay at 9% annually, compounded twice a year. Find the value of your investment in 8 years, expressing the answer in constant pesos.
Answer by lwsshak3(11628)

You wish to invest 2,000 pesos in Uruguay at 9% annually, compounded twice a year. Find the value of your investment in 8 years, expressing the answer in constant pesos.
***
Compound interest formula: A=P(1+i)^n, P=initial investment, i=interest per compounding period, n=number of compounding periods, A=amount after n-periods
..
For given problem:
P=2000
i=.09/2=.045
n=8*2=16
..
A=2000(1+.045)^16
A=2000(1.045)^16
A≈4045
Value of investment in 8 years=4045 pesos